OpenStudy (anonymous):

How do i get the level curves of a function like f(x,y)=(x-y)/(x^2+y^2+1)

OpenStudy (anonymous):

I know that for other problems to find the level curves i let the fuction equal constants. But i just dont know how to start it off when its a fraction like this.

OpenStudy (anonymous):

Its university maths.

OpenStudy (anonymous):

No its calculus i think

OpenStudy (anonymous):

@Tutor.Stacey Are you some computerising tutor or something or some helping robot? In every question thread you say the same thing. "May I get to know the level of math?" "All right I'm working on your question right now".. lol Perhaps you should start working for Tutor.com. When you go on there, you get redirected to this page where this tutor guy starts to persuade you in to buying a lesson and talks just like you =] @zzr0ck3r you agree wid me?

OpenStudy (anonymous):

Its only just the start of it i would need really and then a brief outline?

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

yeah she took off on me

OpenStudy (zzr0ck3r):

now we need to figure out what S.T.A.C.E.Y stands for @genius12

OpenStudy (zzr0ck3r):

we are playing sorry

OpenStudy (zzr0ck3r):

joking

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

anyone have any ideas?

OpenStudy (anonymous):

relax lol.

OpenStudy (anonymous):

We may have gotten off topic a small bit??!

OpenStudy (anonymous):

even if someone knew how to get the gradient of something like this. I know that i get the partials of the equation but when i get them they seem to be wrong?

OpenStudy (anonymous):

its the fraction thats putting me off!

OpenStudy (anonymous):

can anyone answer mine?!

zepdrix (zepdrix):

Hmm, ok I've got an idea.

zepdrix (zepdrix):

Let's try setting it equal to a constant, like you normally do.\[\large k=\frac{x-y}{x^2+y^2+1}\]Multiplying through by our denominator,\[\large k(x^2+y^2+1)=x-y\]Multiplying some stuff,\[\large kx^2+ky^2+k=x-y\]Rearranging some stuff,\[\large kx^2-x+ky^2+y=-k\]Dividing both sides by k,\[\large x^2-\frac{1}{k}x+y^2+\frac{1}{k}y=-1\]This 1/k is just a constant, let's write it as something simpler.\[\large x^2-cx+y^2+cy=-1\]Completing the square on x and y gives us something like this,\[\large \left(x-\frac{c}{2}\right)^2-\frac{c^2}{4}+\left(y+\frac{c}{2}\right)^2-\frac{c^2}{4}=-1\] Which uhhhhhh simplifies to,\[\large \left(x-\frac{c}{2}\right)^2+\left(y+\frac{c}{2}\right)^2=\frac{c^2}{2}-1\] Lot of messy constants, but it looks like the levels are circles. Lemme graph it a sec, to make sure I didn't do anything horrendously stupid, cause I probably did.

zepdrix (zepdrix):

Hmm yah it doesn't look like circles.. did something stupid. Hmm... back to the drawing board..

zepdrix (zepdrix):

@oldrin.bataku Mr smarty pants can prolly figure this out :) heh

OpenStudy (amistre64):

im thinking ... when x=y \[z=\frac{0}{2y^2+1}\] when y=-x \[z=\frac{2x}{2x^2+1}\] when x=-y \[z=\frac{-2y}{2y^2+1}\] etc ....

OpenStudy (amistre64):

letting y = kx, or having some sort of angle setup is another thought in general